# A Prospect-Theoretic Policy Gradient Algorithm for Behavioral Alignment in Reinforcement Learning

### v2

@@ Abstract
Classical reinforcement learning (RL) typically assumes rational decision-making based on expected utility theory. However, this model has been shown to be empirically inconsistent with actual human preferences, as evidenced in psychology and behavioral economics. Cumulative Prospect Theory (CPT) provides a more nuanced model for human-based decision-making, capturing diverse attitudes and perceptions toward risk, gains, and losses. While prior work has integrated CPT with RL to solve a CPT policy optimization problem, the understanding and practical impact of this formulation remain limited. We revisit the CPT-RL framework, offering new theoretical insights into the nature of optimal policies. We further derive a novel policy gradient theorem for CPT objectives, generalizing the foundational result in standard RL. Building on this theorem, we design a model-free policy gradient algorithm for solving the CPT-RL problem and demonstrate its performance through simulations. Notably, our algorithm scales better to larger state spaces compared to existing zeroth-order methods. This work advances the integration of behavioral decision-making into RL.
@@ 1 Introduction
In classical reinforcement learning (RL), rational agents make decisions to maximize their expected cumulative rewards through interaction with their environment. This paradigm is primarily guided by expected utility theory, which has long been the dominant framework for decision-making under risk. Within this framework, agents are generally considered to be risk-neutral; however, risk-seeking and risk-averse behaviors can also be modeled by modifying the utility function, thereby adjusting the policy optimization objective (see e.g. Prashanth et al. (2022) for a survey).
@@
While risk-sensitive RL extends beyond risk-neutral settings to capture individual risk preferences (using e.g. variance or conditional value at risk metrics), it remains limited in fully representing the complexities of human decision-making and fail to capture the psychological nuances of how humans perceive risks, gains, and losses, often leading to decisions misaligned with human preferences. Specifically, risk-sensitive RL fails to account for the asymmetric perception of gains and losses, as well as the probability distortions inherent in human cognition, such as the tendency to overestimate rare events and underestimate frequent ones (see appendix B for an illustration in a simulation). These psychological and cognitive biases, critical to understanding real-world human decisions under uncertainty, are beyond the scope of traditional risk-sensitive RL frameworks, necessitating a more comprehensive approach.
@@
Cumulative Prospect Theory (CPT), developed by the pioneering behavioral psychologist Daniel Kahneman [^1] Daniel Kahneman has been awarded the Nobel Prize in Economic Sciences in 2002 ‘for having integrated insights from psychological research into economic science, especially concerning human judgment and decision-making under uncertainty’. together with his colleague Amos Tversky in their seminal works combining cognitive psychology and economics (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992), provides a more nuanced understanding of human decision-making to explain several empirical observations invalidating the standard expected utility theory. CPT models how individuals perceive outcomes asymmetrically, being risk-averse in the domain of gains and risk-seeking in the domain of losses, and distort probabilities to reflect cognitive biases. These insights have led to widespread applications of CPT in stateless static settings in high-stakes domains such as healthcare in psychiatry (Sip et al., 2018; George et al., 2019; Mkrtchian et al., 2023), chronic diseases treatment (Zhao et al., 2023) and emergency decision making (Sun et al., 2022) where taking into account psychological factors is of paramount importance, as well as other application domains such as energy (Ebrahimigharehbaghi et al., 2022; Dorahaki et al., 2022) and finance (Ladrón de Guevara Cortés et al., 2023; Luxenberg et al., 2024) to name a few. However, these applications often overlook the sequential decision-making nature inherent in many real-world problems, where the outcome of one decision can affect future choices, a critical aspect of RL.
@@
The integration of CPT into RL provides a promising avenue for behaviorally-aligned sequential decision-making, as it allows RL agents to consider both risk preferences and probability distortions in dynamic environments. While a few isolated recent works have explored this integration (L.A. et al., 2016; Borkar and Chandak, 2021; Ramasubramanian et al., 2021; Danis et al., 2023), the understanding and practical impact of CPT in RL remains limited. Specifically, the computational challenges and theoretical underpinnings of CPT-based RL models are still underexplored, and the practical effectiveness of such models in sequential decision-making problems is not well understood.
@@
In this work, we focus on the policy optimization problem where the objective is the CPT value of the cumulative sum of rewards, induced by a parametrized policy in a Markov Decision Process. Our main contributions are as follows:
About optimal policies in CPT-RL. We provide theoretical insights about the nature of an optimal policy for CPT policy optimization. Unlike in standard MDPs where policies are typically Markovian, an optimal policy is stochastic and non-Markovian in general. When we set the probability distortion function to identity, we show that the policy search set can be significantly reduced to a much smaller policy class when solving (CPT-PO). In this same setting, we also characterize a family of utility functions (affine and exponential utility functions) for which the CPT value objective can be maximized with a Markovian policy. However, we prove that this characterization does not hold anymore when considering nontrivial probability distortion and (nonlinear) utilities together in (CPT-PO).
@@
Policy gradient theorem for CPT-RL. We establish a policy gradient theorem providing a closed form expectation expression for the gradient of our CPT-value objective w.r.t. the policy parameter under suitable regularity conditions on the utility and probability distortion functions. This result generalizes the standard policy gradient theorem in RL.
Policy gradient algorithm for CPT-RL. Building on our policy gradient theorem, we design a policy gradient algorithm to solve the CPT policy optimization problem. The stochastic policy gradient we use involves a novel integral term to be computed and we propose a tailored Monte Carlo estimation procedure to approximate it. We perform simulations to illustrate our theoretical results on simple examples. In particular, we test our PG algorithm in two CPT-RL applications: a traffic control application with finite discrete state action spaces and an electricity management task in a continuous state action setting. We also compare the performance of our PG algorithm to the previously proposed zeroth order algorithm to show the robustness and scalability of our algorithm to higher dimensional MDPs.
2 From Classical RL to CPT-RL
@@
Markov Decision Process. A discrete-time Markov Decision Process (MDP) (Puterman, 2014) is a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\rho)$, where $\mathcal{S},\mathcal{A}$ are respectively the state and action spaces, supposed to be finite for simplicity, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the state transition probability kernel, and $r:\mathcal{S}\times\mathcal{A}\to[-r_{\max},r_{\max}]$ is the reward function which is bounded by $r_{\max}>0$, $\rho$ is the initial state probability distribution. A randomized stationary Markovian policy, which we will simply call a policy, is a mapping $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ which specifies for each $s\in\mathcal{S}$ a probability measure over the set of actions $\mathcal{A}$ by $\pi(\cdot|s)\in\Delta(\mathcal{A})$ where $\Delta(\mathcal{A})$ is the simplex over the finite action space $\mathcal{A}$. Each policy $\pi$ induces a discrete-time Markov reward process $\{(s_{t},r_{t}:=r(s_{t},a_{t}))\}_{t\in\mathbb{N}}$ where $s_{t}\in\mathcal{S}$ represents the state of the system at time $t$ and $r_{t}$ corresponds to the reward received when executing action $a_{t}\in\mathcal{A}$ in state $s_{t}\in\mathcal{S}$. We denote by $\mathbb{P}_{\rho,\pi}$ the probability distribution of the Markov chain $(s_{t},a_{t})_{t\in\mathbb{N}}$ generated by the MDP controlled by the policy $\pi$ with initial state distribution $\rho$. We use $\mathbb{E}_{\rho,\pi}$ (or often simply $\mathbb{E}$ instead) to denote the expectation w.r.t. the distribution of the Markov chain $(s_{t},a_{t})_{t\in\mathbb{N}}$. At each time step $t\geq 0$, the agent follows its policy $\pi$ by selecting an action $a_{t}$ drawn from the action distribution $\pi_{t}(\cdot|s_{t})$ where $s_{t}$ is the environment state at time $t$. Then the environment transitions to a state $s_{t+1}$ sampled from the state distribution $\mathcal{P}(\cdot|s_{t},a_{t})$ given by the state transition kernel $\mathcal{P}$ and the agent obtains a reward $r_{t}$. In traditional RL, the goal of the agent in MDPs is to find a policy $\pi$ maximizing the expected cumulative rewards, i.e. the so-called expected return $J(\pi):=\mathbb{E}_{\rho,\pi}[\sum_{t=0}^{H-1}r_{t}]$ where $s_{0}$ follows the initial state distribution $\rho$ and $H\geq 1$ is a finite horizon.
@@
Policy classes. We now introduce different sets of policies which will be important for stating our results. Each policy class is defined according to the information history the policies have access to for selecting actions. Here, a history $h_{t}\in\mathcal{H}$ is a finite sequence of successive states, actions and rewards: $(s_{0},a_{0},r_{0},....,s_{t-1},a_{t-1},r_{t-1})\,.$ [^2] Rewards can be discarded from the history when they are deterministic functions of state-action pairs. More specifically, throughout this work, we will consider the following sets of policies: $\Pi_{NM}:=\{\mathcal{H}\rightarrow\Delta(\mathcal{A})\}$ is the set of non-(necessarily)Markovian policies, $\Pi_{\Sigma,NS}:=\{\mathcal{S}\times\mathbb{R}\times\mathbb{N}\rightarrow\Delta(\mathcal{A})\}$ is the set of policies that only depend on the current state, the timestep and the sum of rewards accumulated so far (i.e. $\pi(s,\sum_{k=0}^{t-1}r_{k},t)$), $\Pi_{\Sigma,S}:=\{\mathcal{S}\times\mathbb{R}\rightarrow\Delta(\mathcal{A})\}$ is the set of policies that only depend on the state and the sum of rewards (i.e. $\pi(s,\sum_{k=0}^{t-1}r_{k})$), $\Pi_{M,NS}:=\{\mathcal{S}\times\mathbb{N}\rightarrow\Delta(\mathcal{A})\}$ is the set of Markovian policies (i.e. $\pi(s,t)$) and $\Pi_{M,S}:=\{\mathcal{S}\rightarrow\Delta(\mathcal{A})\}$ is the set of stationary Markovian policies, i.e. Markovian policies which are time-independent. Deterministic policies assign a single action to each state. For each set of policies defined above, we define their corresponding subset of deterministic policies: $\Pi_{NM}^{D},\Pi_{\Sigma,NS}^{D},\Pi_{\Sigma,S}^{D},\Pi_{M,NS}^{D}$ and $\Pi_{M,S}^{D}$.
@@
$\Pi_{M,S}\subseteq\Pi_{M,NS}\subseteq\Pi_{\Sigma,NS}\subseteq\Pi_{NM}$ and $\Pi_{M,S}\subseteq\Pi_{\Sigma,S}\subseteq\Pi_{\Sigma,NS}\subseteq\Pi_{NM}\,$ (see Fig. 4, App. A).
Cumulative Prospect Theory Value. Instead of the expected return, CPT prescribes to consider the CPT value which will be defined in this paragraph. As previously mentioned, CPT relies on three distinct elements which we further detail in the following:
(a) A reference point. The human agent has a reference attainable reward value in comparison to which they evaluate their possible reward outcomes. Rewards larger than the reference are perceived as gains whereas lower values are viewed as losses.
(b) A utility function $\mathcal{U}:\mathbb{R}\to\mathbb{R}_{+}$. The agent’s utility is a continuous and non-decreasing function which is not necessarily linear w.r.t. the total reward received by the agent. We consider the function $u^{+}:\mathbb{R}\to\mathbb{R}_{+}$ describing the gains and defined for every $x\in\mathbb{R}$ by $u^{+}(x)=\mathcal{U}(x)$ if $x\geq x_{0}$ and zero otherwise. Similarly, the function $u^{-}:\mathbb{R}\to\mathbb{R}_{+}$ which encodes the losses is defined by $u^{-}(x)=-\mathcal{U}(x)$ if $x\leq x_{0}$ and zero otherwise. Here, $x_{0}$ denotes the reference point. Typically, the utility function is concave (respectively convex) for positive (resp. negative) rewards w.r.t. the reference point, i.e. $u^{+}$ is concave on $\mathbb{R}_{+}$ and $-u^{-}$ is convex on $\mathbb{R}_{-}$. For concreteness, we will use Kahneman and Tversky (1979)’s utility function as a running example: $\mathcal{U}(x)=(x-x_{0})^{\alpha}$ if $x\geq x_{0}$ and $\mathcal{U}(x)=-\lambda(x-x_{0})^{\alpha}$ if $x<x_{0}$, where $\lambda=2.25,\alpha=0.88$ are recommended hyperparameters. See Fig. 8 for an illustration with $x_{0}=0\,.$
(c) A probability distortion function $w:[0,1]\to[0,1]$. This is a continuous non-decreasing weight function that distorts the probability distributions of the gain and loss variables. This function typically captures the human tendency to overestimate the probability of rare events and underestimate the probability of more certain ones. Similarly to the utility function, we denote by $w^{+}$ (resp. $w^{-}$) the function that warps the cumulative distribution function for gains (resp.for losses). Both functions are required to be defined on $[0,1]$, with values in $[0,1]$ and to be non-decreasing, continuous, with $w^{+}(0)=w^{-}(0)=0$ and $w^{-}(1)=w^{-}(1)=1$. Examples of such weights functions in the litterature include $w:p\mapsto p^{\eta}(p^{\eta}+(1-p)^{\eta})^{-\frac{1}{\eta}}$ (Kahneman and Tversky, 1979) and $w:p\mapsto\exp(-(-\ln p)^{\eta})$ (Prelec, 1998) where $\eta\in(0,1)$ is a hyperparameter. We refer the reader to appendix F.2 for examples and plots of utility and probability weight functions.
@@
[…]
where appropriate integrability assumptions are assumed. While the CPT value $\mathbb{C}(X)$ accounts for the human agent’s distortions in perception, it also recovers the expectation $\mathbb{E}(X)$ with weight functions $w^{+},w^{-}$ and utility functions $u^{+}$ (resp. $-u^{-}$) restricted to $\mathbb{R}_{+}$ (resp. $\mathbb{R}_{-}$) are set to be the identity functions. In addition, several risk measures such as variance, Conditional Value at Risk (CVar) and distortion risk measures are also particular cases of CPT values with discontinuous probability weighting functions (at a single point). See appendix F for proofs of these facts and Table 1 therein for a synthetic view of the settings captured by CPT. Note that VaR and CVaR are rooted in the financial risk management literature and rely on objective probability distributions whereas CPT originates from behavioral economics and focuses on subjective risk perception and decision-making.
Problem formulation: CPT-RL. In this work, we will focus on the policy optimization problem where the objective is the CPT value of the random variable $X=\sum_{t=0}^{H-1}r_{t}$ recording the cumulative rewards induced by the MDP and the policy $\pi$ for the finite horizon $H\geq 1$:

### v3

# A Prospect-Theoretic Policy Gradient Framework for Behaviorally Nuanced Reinforcement Learning

@@ Abstract
Classical reinforcement learning (RL) typically assumes rational decision-making based on expected utility theory. However, this model has been shown to be empirically inconsistent with actual human preferences, as evidenced in psychology and behavioral economics. Cumulative Prospect Theory (CPT) provides a more nuanced model for human-based decision-making, capturing diverse attitudes and perceptions toward risk, gains, and losses. While prior work has integrated CPT with RL to solve CPT policy optimization problems, the understanding and impact of this formulation remain limited. Our contributions are as follows: (a) we derive a novel policy gradient theorem for CPT objectives, generalizing the foundational result in standard RL, (b) we design a model-free policy gradient algorithm for solving the CPT-RL problem, (c) we analyze our policy gradient estimator and prove asymptotic convergence of the algorithm to first-order stationary points, and (d) test its performance through simulations. Notably, our first-order policy gradient algorithm scales better than existing zeroth-order methods to larger state spaces. Our theoretical framework offers more flexibility to advance the integration of behavioral decision-making into RL.
@@ 1 Introduction
In classical reinforcement learning (RL), rational agents make decisions to maximize their expected cumulative rewards through interaction with their environment. This paradigm is primarily guided by expected utility theory, which has long been the dominant framework for decision-making under risk. Within this framework, agents are generally considered to be risk-neutral; however, risk-seeking and risk-averse behaviors can also be modeled by modifying the utility function, thereby adjusting the policy optimization objective (see e.g. Prashanth et al. (2022); Biswas and Borkar (2023); Bäuerle and Jaśkiewicz (2024) for recent surveys).
@@
While risk-sensitive RL extends beyond risk-neutral settings to capture individual risk preferences (using e.g. variance or conditional value at risk metrics), it remains limited in fully representing the complexities of human decision-making and fail to capture the psychological nuances of how humans perceive risks, gains, and losses, often leading to decisions misaligned with human preferences. Specifically, risk-sensitive RL fails to account for the asymmetric perception of gains and losses, as well as the probability distortions inherent in human cognition, such as the tendency to overestimate rare events and underestimate frequent ones (see App. C for an illustration in a simulation). These psychological and cognitive biases, critical to understanding real-world human decisions under uncertainty, are beyond the scope of traditional risk-sensitive RL frameworks, necessitating a more comprehensive approach.
@@
Cumulative Prospect Theory (CPT), developed by the pioneering behavioral psychologist Daniel Kahneman111was awarded the Nobel Prize in Economic Sciences in 2002 ‘for having integrated insights from psychological research into economic science, especially concerning human judgment and decision-making under uncertainty’. together with his colleague Amos Tversky in their seminal works combining cognitive psychology and economics (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992), provides a more nuanced understanding of human decision-making to explain several empirical observations invalidating the standard expected utility theory. CPT models how individuals perceive outcomes asymmetrically, being risk-averse in the domain of gains and risk-seeking in the domain of losses, and distort probabilities to reflect cognitive biases. These insights have led to widespread applications of CPT in stateless static settings in high-stakes domains such as healthcare in psychiatry Sip et al. (2018); George et al. (2019); Mkrtchian et al. (2023), chronic diseases treatment Zhao et al. (2023) and emergency decision making Sun et al. (2022) where taking into account psychological factors is of paramount importance, as well as other application domains such as energy Ebrahimigharehbaghi et al. (2022); Dorahaki et al. (2022) and finance Ladrón de Guevara Cortés et al. (2023); Luxenberg et al. (2024) to name a few. However, these applications often overlook the sequential decision-making nature of many real-world problems, where the outcome of one decision can affect future choices, a critical aspect of RL.
@@
The integration of CPT into RL provides a promising avenue for behaviorally-aligned sequential decision-making, as it allows RL agents to consider both risk preferences and probability distortions in dynamic environments. While a few isolated recent works have explored this integration (L.A. et al., 2016; Borkar and Chandak, 2021; Ramasubramanian et al., 2021; Danis et al., 2023; Foo et al., 2023), the understanding and practical impact of CPT in RL remains limited. Specifically, the computational challenges and theoretical underpinnings of CPT-based RL models are still underexplored, and the practical effectiveness of such models in sequential decision-making problems is not well understood, especially due to the lack of efficient policy gradient algorithms like in standard RL.
@@
Contributions. In this work, we focus on the policy optimization problem where the objective is the CPT value of the cumulative sum of rewards, induced by a parametrized policy in a Markov Decision Process. Our main contributions are summarized as follows:
@@
(i) Policy gradient theorem for CPT-RL. We establish a policy gradient theorem providing a closed form expectation expression for the gradient of our CPT-value objective w.r.t. the policy parameter under suitable regularity conditions on the utility and probability distortion functions. This result generalizes the standard policy gradient theorem in RL.
(ii) Policy gradient algorithm for CPT-RL. Building on our policy gradient theorem, we design a policy gradient (PG) algorithm to solve the CPT policy optimization problem. To the best of our knowledge, this is the first PG algorithm for CPT-RL using first-order information which does not rely on zeroth-order policy gradient estimation.
(iii) Convergence and sample complexity. We show that our PG estimator is consistent and we analyze its sample complexity. Notably, our sample complexity only scales logarithmically with the dimension of the policy parameter, in contrast to existing zeroth-order PG estimators. We further show that the iterates of our PG algorithm converge asymptotically to the set of stationary points of the CPT-RL objective which is highly non-convex. In particular, unlike prior work, we do not require increasing batch sizes for our PG estimator to guarantee convergence.
(iv) Simulations. We test our PG algorithm on several applications to illustrate the flexibility of CPT-RL compared to standard and risk-sensitive RL in leading to more nuanced behavior. We also compare the performance of our PG algorithm to the previously proposed zeroth order algorithm to show the robustness of our algorithm to larger MDPs.
This paper contributes to the development of more efficient algorithms for human-centered sequential decision-making with models able to incorporate cognitive and psychological biases.
2 From Classical to CPT-RL
2.1 MDPs, CPT Values and Notation
@@
Markov Decision Process. A discrete-time Markov Decision Process (MDP) (Puterman, 2014) is a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,H,\rho)$, where $\mathcal{S},\mathcal{A}$ are respectively the state and action spaces, supposed to be finite for simplicity, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the state transition probability kernel, $r:\mathcal{S}\times\mathcal{A}\to[-r_{\max},r_{\max}]$ is the reward function bounded by $r_{\max}>0$, $\rho$ is the initial state probability distribution and $H\geq 1$ is a finite horizon. A randomized stationary Markovian policy, which we will simply call a policy, is a mapping $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ which specifies for each $s\in\mathcal{S}$ a probability measure over the set of actions $\mathcal{A}$ by $\pi(\cdot|s)\in\Delta(\mathcal{A})$ where $\Delta(\mathcal{A})$ is the simplex over the finite action space $\mathcal{A}$. Each policy $\pi$ induces a discrete-time Markov reward process $\{(s_{t},r_{t}:=r(s_{t},a_{t}))\}_{t\in\mathbb{N}}$ where $s_{t}\in\mathcal{S}$ represents the state of the system at time $t$ and $r_{t}$ corresponds to the reward received when executing action $a_{t}\in\mathcal{A}$ in state $s_{t}\in\mathcal{S}$. We denote by $\mathbb{P}_{\rho,\pi}$ the probability distribution of the Markov chain $(s_{t},a_{t})_{t\in\mathbb{N}}$ generated by the MDP controlled by policy $\pi$ with initial state distribution $\rho$. We use $\mathbb{E}_{\rho,\pi}$ (or often simply $\mathbb{E}$) to denote the expectation. In classical RL, the goal of the agent is to find a policy $\pi$ maximizing the expected return $J(\pi):=\mathbb{E}_{\rho,\pi}[\sum_{t=0}^{H-1}r_{t}]$ where $s_{0}\sim\rho$ and $H\geq 1$ is a finite horizon.
@@
Policy classes. We now introduce different sets of policies which will be important for stating our results. Each policy class is defined according to the information history the policies have access to for selecting actions. Here, a history $h_{t}\in\mathcal{H}:=\bigcup_{h\in[H]}\mathcal{H}_{h}$ (where $\mathcal{H}_{h}:=\mathcal{S}^{h}\times\mathcal{A}^{h}\times[-r_{\max},r_{\max}]^{h}$) is a finite sequence of successive states, actions and rewards: $(s_{0},a_{0},r_{0},...,s_{t-1},a_{t-1},r_{t-1})\,.$ [^2] Rewards can be discarded from the history when they are deterministic functions of state-action pairs. More specifically, throughout this work, we will consider the following sets of policies: $\Pi_{NM}:=\{\mathcal{H}\rightarrow\Delta(\mathcal{A})\}$ is the set of policies that are not necessarily Markovian, $\Pi_{\Sigma,NS}:=\{\mathcal{S}\times\mathbb{R}\times\mathbb{N}\rightarrow\Delta(\mathcal{A})\}$ is the set of non-stationary policies that only depend on the current state, the sum of rewards accumulated so far and the timestep (i.e. $\pi(s,\sum_{k=0}^{t-1}r_{k},t)$), $\Pi_{\Sigma,S}:=\{\mathcal{S}\times\mathbb{R}\rightarrow\Delta(\mathcal{A})\}$ is the set of policies that only depend on the state and the sum of rewards (i.e. $\pi(s,\sum_{k=0}^{t-1}r_{k})$), $\Pi_{M,NS}:=\{\mathcal{S}\times\mathbb{N}\rightarrow\Delta(\mathcal{A})\}$ is the set of (non-stationary) Markovian policies (i.e. $\pi(s,t)$) and $\Pi_{M,S}:=\{\mathcal{S}\rightarrow\Delta(\mathcal{A})\}$ is the set of stationary Markovian policies, i.e. Markovian policies which are time-independent. Deterministic policies assign a single action to each state. For each set of policies defined above, we define their corresponding subset of deterministic policies with a superscript $D$, e.g. $\Pi_{NM}^{D}$.
@@
$\Pi_{M,S}\subseteq\Pi_{M,NS}\subseteq\Pi_{\Sigma,NS}\subseteq\Pi_{NM}$ and $\Pi_{M,S}\subseteq\Pi_{\Sigma,S}\subseteq\Pi_{\Sigma,NS}\subseteq\Pi_{NM}\,$, see Fig. 4 in App. B.
Cumulative Prospect Theory Value. Instead of the expected return, CPT prescribes to consider the CPT value which relies on three distinct elements:
(a) A reference point $x_{0}$. Rewards larger than the reference are perceived as gains whereas lower values are viewed as losses.
(b) A utility function $\mathcal{U}:\mathbb{R}\to\mathbb{R}_{+}$. The agent’s utility is a continuous and non-decreasing function which is not necessarily linear w.r.t. the total reward received by the agent. The function $u^{+}:\mathbb{R}\to\mathbb{R}_{+}$ describing the gains is defined for every $x\in\mathbb{R}$ by $u^{+}(x)=\mathcal{U}(x)$ if $x\geq x_{0}$ and zero otherwise. Similarly, the function $u^{-}:\mathbb{R}\to\mathbb{R}_{-}$ which encodes the losses is defined by $u^{-}(x)=-\mathcal{U}(x)$ if $x\leq x_{0}$ and zero otherwise. Typically, the utility function is concave (resp. convex) for positive (resp. negative) rewards w.r.t. the reference point, i.e. $u^{+}$ is concave on $\mathbb{R}_{+}$ and $u^{-}$ is convex on $\mathbb{R}_{-}$. For concreteness, we will use Kahneman and Tversky (1979)’s utility function as a running example: $\mathcal{U}(x)=(x-x_{0})^{\alpha}$ if $x\geq x_{0}$ and $\mathcal{U}(x)=-\lambda(x-x_{0})^{\alpha}$ if $x<x_{0}$, where $\lambda=2.25,\alpha=0.88$ are recommended hyperparameters. See Fig. 8 for an illustration with $x_{0}=0\,.$
(c) A probability distortion function $w:[0,1]\to[0,1]$. This is a continuous non-decreasing weight function that distorts the probability distributions of the gain and loss variables. This function typically captures the human tendency to overestimate the probability of rare events and underestimate the probability of more certain ones. Similarly to the utility function, we denote by $w^{+}$ (resp. $w^{-}$) the function that warps the cumulative distribution function for gains (resp.for losses). Both functions are required to be defined on $[0,1]$, with values in $[0,1]$ and to be non-decreasing, continuous, with $w^{+}(0)=w^{-}(0)=0$ and $w^{-}(1)=w^{-}(1)=1$. Examples of such weights functions in the literature include $w:p\mapsto p^{\eta}(p^{\eta}+(1-p)^{\eta})^{-\frac{1}{\eta}}$ (Kahneman and Tversky, 1979) and $w:p\mapsto\exp(-(-\ln p)^{\eta})$ (Prelec, 1998) where $\eta\in(0,1)$ is a hyperparameter. We refer the reader to App. G.3 for examples and plots of utility and probability weight functions.
@@
$\mathbb{C}(X)=\int_{0}^{+\infty}w^{+}(\mathbb{P}(u^{+}(X)>z))dz-\int_{0}^{+\infty}w^{-}(\mathbb{P}(u^{-}(X)>z))dz\,$, (1)
where appropriate integrability assumptions are assumed. While the CPT value $\mathbb{C}(X)$ accounts for the human agent’s distortions in perception, it also recovers the expectation $\mathbb{E}(X)$ with weight functions $w^{+},w^{-}$ and utility functions $u^{+}$ (resp. $u^{-}$) restricted to $\mathbb{R}_{+}$ (resp. $\mathbb{R}_{-}$) are set to be the identity functions. In addition, several risk measures such as variance, Conditional Value at Risk (CVaR) and distortion risk measures are also particular cases of CPT values with discontinuous probability weighting functions (see App. G and Table 1 therein).
2.2 CPT-RL Problem Formulation
In this work, we will focus on the CPT Policy Optimization (CPT-PO) problem where the objective is the CPT value of the random variable $X=\sum_{t=0}^{H-1}r_{t}$ recording the cumulative rewards induced by the MDP and the policy $\pi$ for the finite horizon $H\geq 1$: